OpenStudy (anonymous):

Hi everyone! :o) Need a little help here remember a procedure... The problem is: integral (1/(V-IR) d(EYE) I know the integral of 1 over something is ln(something) so I would have ln(V-IR) but the answer is actually ln(V-IR)/-R... I can't remember what I am missing. A chain rule or something? And how to exactly do I implement that rule to get the correct answer? Thanks!

hartnn (hartnn):

\(\int (1/x)dx = \ln x +c \\ but \\ \int (1/[ax+b])dx \ne \ln[ax+b]+c\)

hartnn (hartnn):

the neat way to do this is to substitute the linear expression as u u =V-IR du = ... ?

OpenStudy (anonymous):

lost connection...sorry...working on the du right now

OpenStudy (anonymous):

uhggg...can't belive i forgot this!...need a hint

hartnn (hartnn):

say u = ax+b since a and b are constants du/dx = a*1 +0 = a so, du/dx = a du = adx

hartnn (hartnn):

and i am pretty sure its not "EYE", its I integral (1/(V-IR) d(I)

OpenStudy (anonymous):

i know for example if u=x^3+3x that du would equal 3x^2+3 du i just cant seem to translate that to u=V-IR...would it be du=(1)-(??)

hartnn (hartnn):

we are differentiating w.r.t I so V, R are constants! u= V-IR du/dI = 0 - R*1 makes sense ?

hartnn (hartnn):

dV/dI = 0 because V is a constant

OpenStudy (anonymous):

let me think about it for a second...

hartnn (hartnn):

d/dx (ax) = a dx/dx = a d/dI (-IR) = -R dI/dI = -R

OpenStudy (anonymous):

i am missing something fundamental that used to make me understand this perfectly...i need to back up a second so try and follow what I am gonna ask...

OpenStudy (anonymous):

i am gonna write out each step...just make sure i am correct

OpenStudy (anonymous):

u=x^3+3x.... d/du = d/dx(x^3+3x)... what would be the next step on the left? I know the right would be 3x^2+3

OpenStudy (anonymous):

is writing d/du correct or should it be du/du? uhg

OpenStudy (anonymous):

if i can just figure out the correct thing to write down on the left, then i know i just multiply both sides by du

hartnn (hartnn):

you're differentiating w.r.t "x" right ? so it will be du/dx on left

OpenStudy (anonymous):

ok thank you...that is what i was missing...so one sec then...i think i can finish

hartnn (hartnn):